# Gait Recognition in Large-scale Free Environment via Single LiDAR

### v2

@@ Abstract
Gait features inherently convey unique geometric and dynamic patterns during walking. However, camera-based recognition faces challenges with ambiguities from varying views and lighting. LiDAR’s ability to capture depth makes it pivotal for robotic perception and holds promise for real-world gait recognition. In this paper, based on a single LiDAR, we present the Hierarchical Multi-representation Feature Interaction Network (HMRNet) for robust gait recognition. Prevailing LiDAR-based gait datasets primarily derive from controlled settings with predefined trajectory, remaining a gap with real-world scenarios. To facilitate LiDAR-based gait recognition research, we introduce FreeGait, a comprehensive gait dataset from large-scale, unconstrained settings, enriched with multi-modal and varied 2D/3D data. Notably, our approach achieves state-of-the-art performance on prior dataset (SUSTech1K) and on FreeGait. Code and dataset will be released upon publication of this paper.
Introduction
Gait is a promising biometric feature for human identification, which is hard to disguise and can be easily captured at a long distance without any intrusive interactions with subjects. To explore its huge potential in real-world applications, such as human retrieval, social security, service robots, etc., gait recognition task by identifying a person from sequences of visual data, has been studied for decades (Niyogi and Adelson 1994; Thour and etc. 2021; Kumar and etc 2021).
@@ Introduction
Existing large-scale LiDAR-based gait datasets (Shen et al. 2023) are predominantly collected within controlled lab environments where subjects follow a predetermined routine in a limited space. This lab setting often prompts participants to walk in a manner they believe is expected, potentially skewing their natural gait. Instructions to walk along a straight line, maintain a certain speed, or repeatedly traverse a marked path can render the collected data artificial, contrasting sharply with the spontaneous and varied movements typical in real-world scenarios. This discrepancy creates a gap between lab-collected gait data and the naturalistic human gait in complex real environments.
@@
To bridge this gap in human gait recognition, we introduce FreeGait, a comprehensive dataset captured in open public areas such as subway exits, school gates, and sidewalks (see Figure LABEL:fig:teaser). This dataset encompasses 1,195 subjects of diverse ages and genders, all walking freely in large-scale, unconstrained settings that reflect true pedestrian behavior in both sparse and crowded conditions. Besides 2D silhouettes and 3D point clouds, FreeGait provides diverse data representations, including 2D/3D poses and 3D Mesh &\&& SMPL models, fostering extensive research opportunities.
FreeGait enhances the gait dataset landscape with several novel features:
@@
Prevailing camera-based methods (Chao and etc 2019; Fan, Peng, and etc. 2020; Hou, Cao, and etc 2020; Lin, Zhang, and etc 2021; Liang and etc. 2022; Fan and etc 2023; Teepe and etc 2021) encounters performance bottlenecks due to limited 2D visual ambiguities with view-dependent, illumination-dependent, and depth-missing properties of images. While RGB-D camera-based methods(Hofmann, Geiger, and etc 2014) have been proposed to extract more depth information for gait features, their limited range and inability to function outdoors hinder their applications. In contrast, LiDAR can capture accurate depth information in large-scale indoor and outdoor scenes, unaffected by light conditions. LiDAR point clouds of subjects represent genuine gait-related geometric and dynamic motion attributes, definitely benefiting gait-based identification. A pioneering LiDAR-based gait recognition study, LiDARGait(Shen et al. 2023), employs range view projected from point clouds, demonstrating the capability and effectiveness of LiDAR. However, its dimensionality reduction inevitably diminishes the undistorted geometric and dynamic gait information in the original 3D point clouds.
@@
Related Work
Figure 1: Examples of diverse visual modalities and 2D/3D representations in FreeGait.
@@
Gait Recognition Methods. Previous gait recognition methods (Thour and etc. 2021; Kumar and etc 2021; Chao and etc 2019; Fan, Peng, and etc. 2020; Hou, Cao, and etc 2020; Lin, Zhang, and etc 2021; Huang and etc 2021; Liang and etc. 2022; Fan and etc 2023) primarily use 2D image representation and fall into appearance-based and model-based approaches. The former (Chao and etc 2019; Fan, Peng, and etc. 2020; Hou, Cao, and etc 2020; Lin, Zhang, and etc 2021; Huang and etc 2021; Liang and etc. 2022; Fan and etc 2023) relies on silhouettes from RGB images, making performance dependent on segmentation quality and causing difficulty in identifying humans with changed clothes or cross views. Model-based methods (Teepe and etc 2021; Liao and etc. 2020; Peng et al. 2021) employ skeletons to capture genuine gait characteristics, but are heavily limited by model-based estimation methods. To address depth ambiguity in 2D, some methods (Yu and etc. 2013; Hofmann, Geiger, and etc 2014; Castro et al. 2020; Marín-Jiménez et al. 2021) employ RGB-D cameras for 3D gait features. Yet, depth cameras are limited to indoor scenes with confined sensing range. Recent 3D representation-based approaches(Zheng and etc. 2022) generate 3D meshes from RGB images, but are sensitive to quality and can introduce cumulative errors. (Ahn et al. 2022; Benedek and etc. 2016) explore view-robust gait recognition frameworks based on LiDAR point cloud. However, they all project 3D point cloud into 2D depth map for feature extraction, losing original 3D geometric information and resulting in limited performance. While LiDARGait (Shen et al. 2023) shows promising results by projecting LiDAR point clouds into dense range views, it still lacks essential geometric and dynamic information from raw point clouds. Our approach synergizes range views with 3D point clouds for superior gait recognition.
@@
Gait Recognition Datasets. Existing open gait datasets (Wang et al. 2003; Yu, Tan, and etc 2006; Tan and etc. 2006; Makihara 2010; Iwama et al. 2012; Zhu, Guo, and etc. 2021; Song and etc. 2022; Li and etc. 2023) primarily use silhouettes and are limited to controlled environments, making them unsuitable for real-world applications. GREW(Zhu, Guo, and etc. 2021) addresses this by collecting a dataset in an open area for practical applications. Gait3D(Zheng and etc. 2022) attempts to solve the distortion problem in view-dependent images, but the auxiliary 3D mesh from RGB images still cannot represent the real depth and 3D motion properties. SZTAKI-LGA(Benedek and etc. 2016) collects a LiDAR-based gait dataset, capturing accurate depth information in large-scale scenes, but includes only 28 subjects and limited sequences, insufficient for evaluating learning-based approaches. A recent work introduces a LiDAR-based gait dataset, SUSTech1K(Shen et al. 2023), with 1,050 subjects. However, it is captured in constrained environments with predefined routes, creating a gap compared to real-life scenarios. To advance LiDAR-based gait recognition for real-world applications, we collect a novel dataset with multi-modal visual data and diverse representations in large-scale, unconstrained environments, providing a more accurate and realistic reflection of gait behaviors in real-life situations.
@@
LiDAR-based Applications. LiDAR captures accurate depth information in large-scale scenarios and is unaffected by light conditions, making it popular in autonomous driving and robotics. Numerous LiDAR-based detection and segmentation methods (Zhu and etc. 2021; Zhang and etc. 2020; Alonso et al. 2020; Ning et al. 2023; Wang et al. 2023) have emerged, playing a significant role in 3D perception. Subsequently, numerous human-centric applications(Liu, Rong, and Sheng 2021; Zhao, Ren, and etc 2022; Dai, Lin, and etc 2022; Xu and etc. 2023; Shen et al. 2023), such as pose estimation, motion capture, gait recognition, scene reconstruction, etc., have incorporated LiDAR to expand usage scenarios and improve performance of solutions by utilizing location and geometry features of LiDAR point clouds. We believe that the accurate depth-sensing capability of LiDAR can also definitely benefits in-the-wild gait recognition.
FreeGait Dataset
@@
Table 1: Comparison with public datasets for gait recognition. “Silh”. and “Infr” mean silhouette and infrared, “Crowd” denotes the capture scenes involving multiple persons at the same time in uncontrolled settings, and “D&N” represents day and night during data acquisition. The dataset marked with is no longer available.
@@
CASIA-B (Yu, Tan, and etc 2006) Camera 124 N/A RGB, Silh.
CASIA-C (Tan and etc. 2006) Camera 153 N/A Infrared, Silh.
TUM-GAID (Hofmann, Geiger, and etc 2014) RGB-D & Audio 305 3.6 Audio, Video, Depth
SZTAKI-LGA (Benedek and etc. 2016) LiDAR 28 N/A Point Cloud
OU-MVLP (Takemura, Makihara, and etc. 2018) Camera 10,307 8 Silh.
GREW (Zhu, Guo, and etc. 2021) Camera 26,345 N/A Silh., 2D/3D Pose, Flow
Gait3D (Zheng and etc. 2022) Camera 4,000 N/A Silh., 2D/3D Pose, 3D Mesh&SMPL
CASIA-E (Song and etc. 2022) Camera 1,014 8 Silh., Infr
CCPG (Li and etc. 2023) Camera 200 N/A Silh., RGB
SUSTech1K (Shen et al. 2023) LiDAR & Camera 1,050 12 Silh., RGB, Point Cloud
@@
Most prevailing datasets(Yu, Tan, and etc 2006; Song and etc. 2022; Takemura, Makihara, and etc. 2018) instruct actors to walk on predefined paths in controlled settings. While some in-the-wild gait datasets(Zhu, Guo, and etc. 2021; Zheng and etc. 2022) have been conducted in real-world contexts, they rely solely on cameras and lack 3D human dynamics. Recently, (Shen et al. 2023) claims a LiDAR-based gait dataset, but within constrained settings as shown in Figure 2. In contrast, our FreeGait is a totally free-gait dataset recorded in varied real-world scenarios, from sparse to crowded pedestrian areas. It contains rich data modalities and representations, offering the potential for advanced gait recognition research in practical settings.
Data Acquisition and Statistics
@@
Figure 2: Examples of predifine walking path in constained environments on SUSTech1K (Shen et al. 2023), contravenes the free gait patterns observed in real-world scenarios.
Annotations
During data preprocessing, 2D silhouettes are obtained using a 2D detection method(Ge et al. 2021) and segmentation method(Wang and etc. 2020), while point clouds are detected and cropped using a 3D detector(Cong and etc. 2022). In our dataset, obtaining consecutive frames of personal gait is particularly challenging due to multiple people in one scene. To achieve more accurate tracking performance, we integrate LiDAR and camera information. For point cloud-based 3D tracking, we employ the Hungarian matching algorithm, while initial ID tracking is performed using Byte-Track(Zhang, Sun, and etc 2022) to refine the results. Importantly, to ensure high-quality annotations, we manually correct the tracking ID for crowded scenarios with occlusions.
Characteristics
FreeGait offers special characteristics not available in existing LiDAR-based gait datasets. We highlight three main distinctions below. A detailed comparison with current public datasets can be found in Table 1.
@@
In real-world scenarios, gait serves as a more practical biometric for long-distance person identification than face or fingerprint recognition. Existing gait datasets(Song and etc. 2022; Shen et al. 2023) typically have a capture distance about 12 meters. With the advantages of LiDAR’s long-range sensing, FreeGait is captured in diverse large-scale real-life scenarios, extending the effective human perception distance to approximately 25 meters. This large-scale capture range enhances FreeGait’s applicability in real-world gait applications.
@@
To authentically represent human gait, FreeGait is recorded in real-world settings. Unlike prior datasets that utilize predefined trajectories in controlled environments (Yu, Tan, and etc 2006; Shen et al. 2023), FreeGait’s subjects walk without any constraints, resulting in diverse and practical view variations, as well as more natural challenging factors such as various carrying and dressing, complex and dynamic background clutters, illumination, walking style and etc.
@@
Since FreeGait is captured in natural environments, frequently features multiple persons per scenario, making gait recognition more challenging than in recent SUSTech1K (Shen et al. 2023). Notably, occlusions in crowded settings bring challenges for gait recognition due to incomplete or noisy data. While these crowded conditions pose annotation challenges, they are crucial for evaluating algorithmic robustness and advancing real-world gait-related research and applications. More dataset statistics and examples are detailed in our supplementay.
Privacy Preservation

### v3

@@ Abstract.
Human gait recognition is crucial in multimedia, enabling identification through walking patterns without direct interaction, enhancing the integration across various media forms in real-world applications like smart homes, healthcare and non-intrusive security. LiDAR’s ability to capture depth makes it pivotal for robotic perception and holds promise for real-world gait recognition. In this paper, based on a single LiDAR, we present the Hierarchical Multi-representation Feature Interaction Network (HMRNet) for robust gait recognition. Prevailing LiDAR-based gait datasets primarily derive from controlled settings with predefined trajectory, remaining a gap with real-world scenarios. To facilitate LiDAR-based gait recognition research, we introduce FreeGait, a comprehensive gait dataset from large-scale, unconstrained settings, enriched with multi-modal and varied 2D/3D data. Notably, our approach achieves state-of-the-art performance on prior dataset (SUSTech1K) and on FreeGait. https://4dvlab.github.io/project_page/FreeGait.html
Gait Recogntion, Human-centric Multimedia Understanding, LiDAR point cloud
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681166isbn: 979-8-4007-0686-8/24/10copyright: ccccs: Computing methodologies Biometricsccs: Computing methodologies Activity recognition and understanding
Figure 1. We propose FreeGait, a new LiDAR-based in-the-wild gait dataset under various crowd density and occlusion across different real-life scenes. FreeGait is captured in diverse large-scale real-life scenarios with free trajectory, resulting in various challenges such as 1. occlusions, 2. noise from crowd, 3. noise from carrying objects and etc., as shown in the right part.
[…]
xx.
1. Introduction
Gait is a promising biometric feature for human identification (Chao and etc, 2019; Fan et al., 2020; Hou et al., 2020; Lin et al., 2021; Huang and etc, 2021; Liang and etc., 2022; Fan and etc, 2023) that is difficult to disguise and can be captured from a distance without intrusive interactions. This characteristic is particularly significant for multimedia applications (Wang et al., 2023a, b; Yu et al., 2022; Zheng et al., 2022; Lin et al., 2020), where the seamless integration of information across various media forms is paramount. For example, in smart homes and IoT environments, gait recognition tailors user experiences by automatically adjusting environmental controls to individual preferences based on their movement patterns. Gait recognition also provides advanced diagnostic and monitoring tools in healthcare, offering insights into patient health and flagging potential issues, thus integrating essential health data with multimedia systems for improved patient management. Additionally, its non-intrusive nature makes gait recognition ideal for security and surveillance, enabling passive monitoring in environments where conventional methods may be too invasive. Despite being a subject of extensive research for decades, human gait recognition continues to be highly relevant and holds significant potential in the evolving multimedia landscape.
@@ 1. Introduction
Existing large-scale LiDAR-based gait datasets (Shen et al., 2023) are predominantly collected within controlled lab environments where subjects follow a predetermined routine in a limited space. This lab setting often prompts participants to walk in a manner they believe is expected, potentially skewing their natural gait. Instructions to walk along a straight line, maintain a certain speed, or repeatedly traverse a marked path can render the collected data artificial, contrasting sharply with the spontaneous and varied movements typical in real-world scenarios. This discrepancy creates a gap between lab-collected gait data and the naturalistic human gait in complex real environments.
@@
To bridge this gap in human gait recognition, we introduce FreeGait, a comprehensive dataset captured in open public areas such as subway exits, school gates, and sidewalks (see Figure 1). This dataset encompasses 1,195 subjects of diverse ages and genders, all walking freely in large-scale, unconstrained settings that reflect true pedestrian behavior in both sparse and crowded conditions. Besides 2D silhouettes and 3D point clouds, FreeGait provides diverse data representations, including 2D/3D poses and 3D Mesh &\&& SMPL models, fostering extensive research opportunities. FreeGait enhances the gait dataset landscape with several novel features:
@@
Prevailing camera-based methods (Chao and etc, 2019; Fan et al., 2020; Hou et al., 2020; Lin et al., 2021; Liang and etc., 2022; Fan and etc, 2023; Teepe and etc, 2021) encounters performance bottlenecks due to limited 2D visual ambiguities with view-dependent, illumination-dependent, and depth-missing properties of images. While RGB-D camera-based methods(Hofmann et al., 2014) have been proposed to extract more depth information for gait features, their limited range and inability to function outdoors hinder their applications. In contrast, LiDAR can capture accurate depth information in large-scale indoor and outdoor scenes, unaffected by light conditions. LiDAR point clouds of subjects represent genuine gait-related geometric and dynamic motion attributes, definitely benefiting gait-based identification. A pioneering LiDAR-based gait recognition study, LiDARGait(Shen et al., 2023), employs range view projected from point clouds, demonstrating the capability and effectiveness of LiDAR. However, its dimensionality reduction inevitably diminishes the undistorted geometric and dynamic gait information in the original 3D point clouds.
@@
2. Related Work
Figure 2. Examples of diverse visual modalities and 2D/3D representations in FreeGait.
@@ 2. Related Work
Gait Recognition Methods. Previous gait recognition methods (Thour and etc., 2021; Kumar and etc, 2021; Wang et al., 2023a, b; Yu et al., 2022; Zheng et al., 2022; Lin et al., 2020) primarily use 2D image representation and fall into appearance-based and model-based approaches. The former (Chao and etc, 2019; Fan et al., 2020; Hou et al., 2020; Lin et al., 2021; Huang and etc, 2021; Liang and etc., 2022; Fan and etc, 2023) relies on silhouettes from RGB images, making performance dependent on segmentation quality and causing difficulty in identifying humans with changed clothes or cross views. Model-based methods (Teepe and etc, 2021; Liao and etc., 2020; Peng et al., 2021) employ skeletons to capture genuine gait characteristics, but are heavily limited by model-based estimation methods. To address depth ambiguity in 2D, some methods (Yu and etc., 2013; Hofmann et al., 2014; Castro et al., 2020; Marín-Jiménez et al., 2021) employ RGB-D cameras for 3D gait features. Yet, depth cameras are limited to indoor scenes with confined sensing range. Recent 3D representation-based approaches(Zheng and etc., 2022) generate 3D meshes from RGB images, but are sensitive to quality and can introduce cumulative errors. (Ahn et al., 2022; Benedek and etc., 2016) explore view-robust gait recognition frameworks based on LiDAR point cloud. However, they all project 3D point cloud into 2D depth map for feature extraction, losing original 3D geometric information and resulting in limited performance. While LiDARGait (Shen et al., 2023) shows promising results by projecting LiDAR point clouds into dense range views, it still lacks essential geometric and dynamic information from raw point clouds. Our approach synergizes range views with 3D point clouds for superior gait recognition.
@@
Gait Recognition Datasets. Existing open gait datasets (Wang et al., 2003; Yu et al., 2006; Tan and etc., 2006; Makihara, 2010; Iwama et al., 2012; Zhu et al., 2021; Song and etc., 2022; Li and etc., 2023) primarily use silhouettes and are limited to controlled environments, making them unsuitable for real-world applications. GREW(Zhu et al., 2021) addresses this by collecting a dataset in an open area for practical applications. Gait3D(Zheng and etc., 2022) attempts to solve the distortion problem in view-dependent images, but the auxiliary 3D mesh from RGB images still cannot represent the real depth and 3D motion properties. SZTAKI-LGA(Benedek and etc., 2016) collects a LiDAR-based gait dataset, capturing accurate depth information in large-scale scenes, but includes only 28 subjects and limited sequences, insufficient for evaluating learning-based approaches. A recent work introduces a LiDAR-based gait dataset, SUSTech1K(Shen et al., 2023), with 1,050 subjects. However, it is captured in constrained environments with predefined routes, creating a gap compared to real-life scenarios. To advance LiDAR-based gait recognition for real-world applications, we collect a novel dataset with multi-modal visual data and diverse representations in large-scale, unconstrained environments, providing a more accurate and realistic reflection of gait behaviors in real-life situations.
@@
LiDAR-based Applications. LiDAR captures accurate depth information in large-scale scenarios and is unaffected by light conditions, making it popular in autonomous driving and robotics. Numerous LiDAR-based detection and segmentation methods (Ning et al., 2023; Liu et al., 2023; Wang et al., 2023c; Chen et al., 2023; Lu et al., 2023; Peng et al., 2023a, b; Yao et al., 2024) have emerged, playing a significant role in 3D perception. Subsequently, numerous human-centric applications(Liu et al., 2021; Zhao et al., 2022; Dai et al., 2022; Xu et al., 2023; Shen et al., 2023; Cong et al., 2024; Ren et al., 2024; Xu et al., 2024), such as pose estimation, motion capture, gait recognition, scene reconstruction, etc., have incorporated LiDAR to expand usage scenarios and improve performance of solutions by utilizing location and geometry features of LiDAR point clouds. We believe that the accurate depth-sensing capability of LiDAR can aslo denefinitly benefits in-the-wild gait recognition.
3. FreeGait Dataset
@@
Table 1. Comparison with public datasets for gait recognition. “Silh”. and “Infr” mean silhouette and infrared, “Crowd” denotes the capture scenes involving multiple persons at the same time in uncontrolled settings, and “D&N” represents day and night during data acquisition. The dataset marked with is no longer available.
@@
CASIA-B (Yu et al., 2006) Camera 124 N/A RGB, Silh.
CASIA-C (Tan and etc., 2006) Camera 153 N/A Infrared, Silh.
TUM-GAID (Hofmann et al., 2014) RGB-D & Audio 305 3.6 Audio, Video, Depth
SZTAKI-LGA (Benedek and etc., 2016) LiDAR 28 N/A Point Cloud
OU-MVLP (Takemura et al., 2018) Camera 10,307 8 Silh.
GREW (Zhu et al., 2021) Camera 26,345 N/A Silh., 2D/3D Pose, Flow
Gait3D (Zheng and etc., 2022) Camera 4,000 N/A Silh., 2D/3D Pose, 3D Mesh&SMPL
CASIA-E (Song and etc., 2022) Camera 1,014 8 Silh., Infr
CCPG (Li and etc., 2023) Camera 200 N/A Silh., RGB
SUSTech1K (Shen et al., 2023) LiDAR & Camera 1,050 12 Silh., RGB, Point Cloud
@@
Most prevailing datasets(Yu et al., 2006; Song and etc., 2022; Takemura et al., 2018) instruct actors to walk on predefined paths in controlled settings. While some in-the-wild gait datasets(Zhu et al., 2021; Zheng and etc., 2022) have been conducted in real-world contexts, they rely solely on cameras and lack 3D human dynamics. Recently, (Shen et al., 2023) claims a LiDAR-based gait dataset, but within constrained settings as shown in Figure 3. In contrast, our FreeGait is a totally free-gait dataset recorded in varied real-world scenarios, from sparse to crowded pedestrian areas. It contains rich data modalities and representations, offering the potential for advanced gait recognition research in practical settings.
3.1. Data Acquisition and Statistics
@@
Figure 3. Examples of predefined walking path in constrained environments on SUSTech1K (Shen et al., 2023), contravenes the free gait patterns observed in real-world scenarios.
3.2. Annotations
During data preprocessing, 2D silhouettes are obtained using a 2D detection method(Ge et al., 2021) and segmentation method(Wang and etc., 2020), while point clouds are detected and cropped using a 3D detector(Cong and etc., 2022). In our dataset, obtaining consecutive frames of personal gait is particularly challenging due to multiple people in one scene. To achieve more accurate tracking performance, we integrate LiDAR and camera information. For point cloud-based 3D tracking, we employ the Hungarian matching algorithm, while initial ID tracking is performed using Byte-Track(Zhang et al., 2022) to refine the results. Importantly, to ensure high-quality annotations, we manually correct the tracking ID for crowded scenarios with occlusions.
3.3. Characteristics
FreeGait offers special characteristics not available in existing LiDAR-based gait datasets. We highlight three main distinctions below. A detailed comparison with current public datasets is shown in Table 1.
@@ 3.3. Characteristics
In real-world scenarios, gait serves as a more practical biometric for long-distance person identification than face or fingerprint recognition. Existing gait datasets(Song and etc., 2022; Shen et al., 2023) typically have a capture distance about 12 meters. With the advantages of LiDAR’s long-range sensing, FreeGait is captured in diverse large-scale real-life scenarios, extending the effective human perception distance to approximately 25 meters. This large-scale capture range enhances FreeGait’s applicability in real-world gait applications.
@@
To authentically represent human gait, FreeGait is recorded in real-world settings. Unlike prior datasets that utilize predefined trajectories in controlled environments (Yu et al., 2006; Shen et al., 2023), FreeGait’s subjects walk without any constraints, resulting in diverse and practical view variations, as well as more natural challenging factors such as various carrying and dressing, complex and dynamic background clutters, illumination, walking style and etc.
@@
Since FreeGait is captured in natural environments, frequently features multiple persons per scenario, making gait recognition more challenging than in recent SUSTech1K (Shen et al., 2023). Notably, occlusions in crowded settings bring challenges for gait recognition due to incomplete or noisy data. While these crowded conditions pose annotation challenges, they are crucial for evaluating algorithmic robustness and advancing real-world gait-related research and applications. More dataset statistics and examples are detailed in our supplementay.
3.4. Privacy Preservation